Grade 11Trigonometry

sinB÷sin(2A+B)=m÷n,prove that cot(A+B)=m-n÷n+m×cotA

Ramanpreet
7 Years agoGrade 11

1 Answer

Aditya Gupta
7 Years ago
sinB/sin(2A+B)=m/n
now apply componendo and dividendo
(sinB – sin(2A+B))/(sinB+sin(2A+B)) = (m-n)/(m+n)
now apply the formulas of sinx+siny and sinx – siny to get
2sin(-A)cos(A+B)/2sin(A+B)cos(A)= (m-n)/(m+n)
 – cot(A+B)/cotA=  (m-n)/(m+n) and hence the result. please approve :)